# Deep Learning based Visually Rich Document Content Understanding: A Survey

### v1

@@ 4.2. Coarse and Joint-grained Pretrained Models
4.2.1. Coarse-grained Frameworks
@@
SelfDoc (Li et al., 2021b) is the first pretrained VRDU model to leverage coarse-grained document elements for various document understanding tasks. Unlike fine-grained models that rely on OCR-extracted text sequences, SelfDoc uses Faster-RCNN to extract Regions of Interest (RoIs) from document entities. This approach reduces the input sequence length for text-dense and long documents with improved time and space complexities. To acquire the initial multimodal representations, Faster-RCNN extracts visual embeddings from these RoIs, while Sentence-BERT (Reimers and Gurevych, 2019) provides textual embeddings based on the OCR-extracted text of each entity. The textual and visual representations are fed into two single-modality BERT-like encoders to learn the intra-modality correlations between entities contextually. A cross-modality encoder is introduced to enhance inter-modality learning, mirroring the structure of single-modality encoders but incorporating cross-attention layers. It randomly masks entities in document images within either language or vision branches during pretraining. Furthermore, a modality-adaptive attention mechanism dynamically adjusts weights for visual and textual embeddings based on the input and task, creating robust entity representations for fine-tuning.
UniDoc (Gu et al., 2021) is an entity-level document understanding model which contains a trainable image encoder with RoI-Align (He et al., 2017) to extract entity visual representation and novel cross-attention mechanisms to fuse multimodal information. The initial sentence and textual embeddings are acquired by RoI features and averaged word embeddings summing up with the linear projected bbox coordinates. To enable the trainable visual representation to predict meaningful visual cues effectively, product quantization (Jegou et al., 2010) is applied to discretise a RoI feature into a finite set of visual representations and mapped to a new embedding. Additionally, a novel Gated Cross-Attention is introduced to improve the interactive learning between various modalities. After a multi-head cross-attention to get the enhanced visual and textual representations, a gating mechanism is applied to dynamically weight the visual ($V$) and textual features ($T$) by feeding concatenated $[V:T]$ to a non-linear network to acquire a modality-aware attention bias $\beta_{v},\beta_{t}$ for visual and textual respectively. There are three pretraining tasks are conducted simultaneously to enhance multimodal feature representatives. Masked Sentence Modelling and Visual Contrastive Learning are required to predict the masked sentence representation and quantize visual representation, respectively. Another Vision-Language Alignment introduced by LayoutLMv2 (Xu et al., 2020b) to enforce contextual learning between multi-modalities. But instead of using split image regions, UniDoc aligns the image and text to belonging to the same entity.
4.2.2. Joint-grained Frameworks
StrucText (Li et al., 2021c) is a multimodal pretrained VRDU model using fine-grained textual and coarse-grained vision information to capture richer geometric and semantic information from different levels and modalities. The layout embedding of each text token and segment is encoded by $L=W_{l}[x0,y0,x1,y1,w,h],W_{1}\in\mathbb{R}^{6\times N}$. They follow token-level models to encode the sequence of tokens and use pretrained ResNet50 with FPN to extract the entity visual features. A segment-ID embedding is allocated to each token’s textual and entity visual representations to boost the alignment learning. Three self-supervised tasks are introduced to enhance inter-modality learning: MVLM, Segment length Prediction (SLP), and Paired Box Direction (PBD). As a new self-supervised learning task, SLP asks the model to predict the entity’s length to leverage the entity’s visual embeddings and textual information from the same segment ID. Another self-supervised learning task aims to learn more comprehensive pair-wise spatial correlations between entities by clarifying their spatial correlation.
Fast-StructText (Zhai et al., 2023) is built on StrucText to improve model efficiency and enhance the feature expressiveness by introducing an hourglass transformer and Symmetry Cross-Attention mechanisms (SCA). Similar feature encoding methods to StructText are adopted, but the bbox formats of layout encoding are $[x0,y0,x1,y1]$. To encode the module and progressively reduce the redundancy tokens, an hourglass transformer is proposed consisting of several Merging blocks and Extension blocks to down-sampling and up-sampling the number of input tokens. Merging blocks suggest merging neighbouring $k$ tokens with the weighted 1D average pooling to shorten the sequence length. The extension block is required to transform the shortened sequence back to the entire length by simply applying a repeat up-sampling method. In addition, to enhance the interaction between textual and vision modalities, SCA consists of two dual cross-attention to handle text and visual features. Different self-supervised tasks are conducted simultaneously, including MVLM (Xu et al., 2020b; Li et al., 2021c), Graph-based Token Relation (GTR), Sentence Order Prediction (SOP) and Text-Image Alignment (Huang et al., 2022). For the two new proposed tasks, GTR is a task similar to Paired-Box Direction to predict the positional relations between paired entity visual features and SOP is used to predict whether two sentence pairs are normal-order adjacent to learning semantic knowledge.
@@ 4.2.2. Joint-grained Frameworks
MGDoc (Wang et al., 2022a) is the first multimodal multi-granular pretrained framework which introduces multi-granular attention and cross-modal attention to increase the inter-grained learning and better cross-modality fusion. For acquiring initial features of different modalities, pretrained language models (Reimers and Gurevych, 2019) and vision encoders (He et al., 2016) are used to encode different-grained inputs from word to whole page. The encoded modalities are associated with positional features (Zhai et al., 2023) and modality-type embeddings (Li et al., 2021c) to acquire the final input representation of various modalities. To encode the hierarchical relation between multi-grained features, including pages, entities and words, two attention biases are added to the original self-attention weights. A hierarchical bias is a binary value of 0 or 1 to examine the inside or outside relation between two inputs, while the relation bias is the relative positive between two bboxes. Apart from multi-grained learning, cross-attention is applied to fuse cross-modality information better. Three tasks are proposed during pretraining, including Mask Text Modeling (MTM), Mask Vision Modeling (MVM) and Multi-Granularity Modeling. MTM and MVM randomly mask the multi-grained textual and visual representations to predict the masked features under the mean absolute error. To boost the interactive understanding of multi-granularity, a token-entity linking prediction is proposed by computing the dot-product between token and entity features.
@@
WUKONG-READER (Bai et al., 2022) uses fine-grained level inputs but leverages coarse-grained level self-supervised tasks to enhance the fine-grained level information. The model inputs contain a document image and OCR extracted sequence of tokens with their bboxes. A Mask-RCNN is used as a visual backbone to acquire several visual tokens and extract visual features of textlines from a RoIHead layer. For textual information encoding, the first is layers of RoBERTa(Wei et al., 2020) are applied to extract the token representation and sum up with additional features following (Xu et al., 2020b). The concatenated visual and textual features are fed into a multimodal modal encoder based on the rest six layers of RoBERTa. Different pretraining tasks are proposed, including MLM, Textline-Region Contrastive Learning (TRC), Masked Region Modeling (MRM) and Textline Grid Matching (TCM). TRC aims to enhance cross-modality interactive learning at the entity (textline) level by following a text-vision-aligned contrastive learning approach (Yao et al., 2021). To enhance the visual representation, MRM is applied to mask 15% of textlines randomly and to predict the masked visual embeddings following (Li et al., 2021b). A text-grid alignment task is introduced to enhance layout understanding by dividing the document image into N-grids and predicting the tokens in 15% selected textlines belonging to which grid. The scaling parameters are applied to each loss of pretraining tasks.
@@
GeoLayoutLM (Luo et al., 2023) is a sophisticated multimodal framework that distinctively incorporates geometric information through specialised pretraining tasks and the development of innovative relation heads. Inspired by the dual-stream structure of METER and SelfDoc (Li et al., 2021b), GeoLayoutLM features separate vision and text-layout modules coupled with interactive co-attention layers that enhance the integration of visual and textual data. The model introduces two advanced relation heads—the Coarse Relation Prediction (CRP) head and the Relation Feature Enhancement (RFE) head—which refine relation feature representation crucial for both pretraining and fine-tuning phases. The pretraining regimen includes tasks designed to understand geometric relationships, such as GeoPair, GeoMPair, and GeoTriplet, aiding the model in grasping the complex dynamics of document layouts. During fine-tuning, the model utilises pretrained parameters to optimise both semantic entity recognition and relation extraction tasks, employing a novel inference technique that enhances relation pair selection accuracy by focusing on the most probable relationships and minimizing variance among potential options.
4.3. Encoder-Decoder Pretrained Frameworks
@@
In addition to the above encoder-only frameworks, researchers have proposed encoder-decoder pretrained models that often approach tasks like Key Information Extraction (KIE) or Visual Question Answering (VQA) in a generative style. Addressing limitations of OCR-dependent frameworks, such as accumulated OCR errors and incorrect reading orders, OCR-independent models have been introduced for end-to-end VRDU.
4.3.1. OCR-dependent Encoder-Decoder Frameworks
@@
TILT (Powalski et al., 2021) is a T5-based transformer encoder-decoder architecture enhanced with a relative spatial bias in the self-attention mechanism to acquire fine-grained token representations. It incorporates encoding methods that apply relative sequence input bias and capture horizontal and vertical distance biases in the attention scores. A U-Net-based framework is applied to extract the fixed-size feature maps fed into the encoder together. They follow T5 pretraining strategies on RVL-CIDP dataset (Harley et al., 2015) but use a salient span masking scheme adopted by (Raffel et al., 2020; Guu et al., 2020). As the first encoder-decoder framework, TILT requires off-the-shelf OCR tools to acquire the textual token sequence. To conduct VRD content understanding end-to-end, some OCR-free frameworks are proposed to solve the limitations of OCR-dependent models.
UDOP (Tang et al., 2022) proposes an encoder-decoder pre-trained document understanding framework that leverages a ViT-based model (Dosovitskiy et al., 2020), following the principles of LayoutLMv3 (Huang et al., 2022), to process multimodal information. To enhance the comprehensiveness of textual token embeddings, the framework sums the text token embeddings with token-aligned image patch embeddings when the centre of the bounding box falls within the image patch. Additionally, the positional biases introduced by TILT are added, but ID positional encoding is not used. The decoding stage is designed to generate all vision, text and layout modalities consisting of a bi-directional Transformer text-layout decoder and an MAE vision decoder. Two decoders are cross-attended with each other. The pertaining tasks contain self-supervised on unlabeled documents to learn robust document representation and supervised pertaining tasks for fine-grained model supervision. Masked text layout and image reconstruction are used to predict the masked information using unmasked multimodal information. A layout moulding task is applied to predict the positions of groups of text tokens, and a visual text recognition task is proposed to identify text at the given location in the image. Supervised tasks utilise a training set of publicly available benchmark datasets of different downstream tasks, including document classification (Harley et al., 2015), KIE (Stanisławek et al., 2021), VQA (Tanaka et al., 2021; Mathew et al., 2021) and layout analysing (Zhong et al., 2019).
@@
DocFormerv2 (Appalaraju et al., 2024) is an encoder-decoder transformer architecture that uses multimodal (visual, textual and positional) to enhance the multimodal understanding and layout-aware language decoder to predict the predictions. The patched image pixels are fed into the convolutional and linear layers to get down-sampled patch embedding. The textual embeddings are acquired by linear projected token one-hot encoding. Both visual and textual embeddings are summed with the 2D-positional encoding (Xu et al., 2020a) of Patches and linear project bbox embedding ($[x0,y0,x1,y1]$) (Wang et al., 2022b), respectively. Two encoder-based Token to Line (T2L), Token to Grid (T2G), and one decoder-based self-supervised learning task, MLM, are proposed to enable multimodal feature interactive learning. T2L aims to improve the relative position understanding between tokens by predicting the number of textlines between two randomly selected tokens. For improving the layout and structure, understanding needs to split the image into $m\times n$ grids to predict the located grid number of each OCR token. For the decoder MLM, the spatial feature of each masked text token is masked as well, and the other setup follows T5 (Raffel et al., 2020).
@@
ViTLP (Mao et al., 2024) introduces an encoder-decoder architecture for OCR and document understanding using a ViT-based vision encoder to obtain image patch representations, which are fed into a decoder to generate text and layout sequences autoregressively. A special ”[LOC]” token, encoding bounding box coordinates $[x0,y0,x1,y1]$, reduces the layout token sequence length. To control generation flow, ”[BOS]” and ”[CONT]” tokens are added, representing input sequences of two tokens, t1,t2, as ”[BOS], t1, [LOC], t2, [LOC]”. The decoder has hierarchical heads: the text head uses all tokens to generate the next text token, while the layout head uses ”[LOC]” tokens to predict bounding box coordinates. The ”[CONT]” token handles sequences of arbitrary length by continuing generation until ”[END]”, based on the prefix token ratio.
4.3.2. OCR-free Pretrained Frameworks
@@
ReRum (Cao et al., 2023a) introduce an end-to-end architecture in which the decoding process focuses on local interest and visual cues. Like other OCR-free frameworks, a Swin-Transformer extracts image patch representations. The extracted visual features are fed into a transformer-based Query-Decoder to acquire the vision-enhanced query representation. The enhanced query representation with visual features is fed into a Text Decoder to generate the predicted text tokens auto-regressively. Notably, a Content-aware Token Merge technique is proposed to dynamically focus on more relevant foreground parts of visual features, which selects top-K visual tokens based on the averaged correlation scores between query and visual token representations. Additionally, the unselected visual tokens (called Background Area) contain rich global features that could be used to enhance the Top-K foreground visual tokens through basic attention mechanisms. Three pretraining tasks are applied: query to segmentation (Q2S), text to segmentation (T2S), and segmentation to text (S2T). Q2S follows DETR setup for a token generation but alters to an instance segmentation task to predict $N$ mask for the target text area to improve text detection ability. T2S acquire the Text Decoder output to conduct another instance segmentation task after feeding the image text snippet into the Text Decoder to boost the layout-aware textual information understanding. The last S2T is to use the outputs from the Text Decoder to generate the token like an OCR task auto-regressively.
@@
StrucTextV2 (Yu et al., 2022) is an end-to-end structure that uses image-only input to conduct several downstream tasks. It contains a CNN-based visual extractor with FPN strategies (Lin et al., 2017) and follows ViT (Dosovitskiy et al., 2020) to get linear projected flattened patch-level representations. The patch token embeddings serve as the input to the Transformer encoder to enhance the contextually semantic representations. Then, the lightweight fusion network is applied to generate the final representations and fed into two branches during pretraining: made language Modeling (MLM) and Masked Image Modeling (MIM). Instead of using text inputs when MLM is used by other models (Devlin, 2018), a portion of the text regions are masked with RGB values $[255,255,255]$ randomly with a 2-layer MLP decoder to predict the masked token. MIM masks the rectangular text regions and predicts the RGB values of the missing pixels to improve the document representations. Except for the global average pooled FPN fused visual representations, the MLM-generated hidden state of each text region is concatenated and fed into a Fully Convolutional New York to get the regressed masked missing pixel values.
4.3.3. LLM-based Frameworks
HRVDA (Liu et al., 2024b) aims to propose a MLLM accepting high-resolution image inputs to conduct fine-grained information extraction from VRDs. A swin-transformer (Liu et al., 2021) is used to encode document images into image patch tokens. A pluggable content detector then identifies visual tokens that contain relevant document content information. Following this, a content filtering mechanism performs token pruning to remove irrelevant tokens. The remaining encoded visual tokens are processed through an MLP to ensure consistency with the LLM embedding space dimensions. These pruned tokens are then fused with instruction features, allowing further filtering of tokens irrelevant to the instructions. The final streamlined set of visual tokens and instructions is fed into the LLM, which generates the corresponding responses.

### v2

@@ 4.2 Coarse and Joint-grained Pretrained Models
4.2.1 Coarse-grained Frameworks
@@
SelfDoc [selfdoc] is the first pretrained VRD-CU model leveraging coarse-grained document semantic entity for various understanding tasks. Unlike fine-grained OCR-based models, it uses Faster-RCNN to extract Regions of Interest (RoIs) from semantic entities, reducing input length and improving efficiency on dense, long documents. Visual embeddings are derived from RoIs, while Sentence-BERT [sentencebert] generates textual embeddings from the OCR-extracted text. These are fed into separate BERT-style encoders to model intra-modality context. A cross-modality encoder with cross-attention layers enables inter-modality learning. During pretraining, random masking is applied to either text or visual tokens. A modality-adaptive attention mechanism further adjusts the weight of visual and textual cues dynamically for robust entity-level representations.
UniDoc [udoc] is an coarse-grained model with a trainable image encoder using RoI-Align [maskrcnn] to extract visual features and a novel cross-attention mechanism for multimodal fusion. It computes textual embeddings by combining average word embeddings and linearly projected bounding box coordinates. Product quantization [jegou2010product] is used to discretize RoI features into a finite visual codebook. A Gated Cross-Attention module is applied after multi-head fusion, where concatenated $[V:T]$ features are passed through a non-linear layer to generate modality-aware attention biases ($\beta_{v}$, $\beta_{t}$). UniDoc employs three pretraining tasks: Masked Sentence Modelling, Visual Contrastive Learning, and Vision-Language Alignment (adapted from LayoutLMv2 [layoutlmv2])—the last aligning image-text pairs at the entity level instead of region-level.
4.2.2 Joint-grained Frameworks
StrucText [structext] is a multimodal pretrained VRD-CU model that integrates fine-grained text and coarse-grained visual features to capture rich geometric and semantic information. Each text token is assigned a layout embedding $L=W_{l}[x_{0},y_{0},x_{1},y_{1},w,h]$ using $W_{l}\in\mathbb{R}^{6\times N}$, and visual features are extracted from entities via a pretrained ResNet50-FPN. A segment-ID embedding aligns text and visual representations within the same entity. Three self-supervised tasks guide training: Masked Visual-Language Modeling (MVLM), Segment Length Prediction (SLP), and Paired Box Direction (PBD). SLP predicts entity lengths to encourage multimodal fusion, while PBD captures spatial relationships between entities.
Fast-StructText [faststructext] builds upon StrucText to improve efficiency and representation power. It uses a more compact layout encoding $[x_{0},y_{0},x_{1},y_{1}]$, an hourglass transformer with Merging and Extension blocks for token compression and recovery, and a Symmetry Cross-Attention (SCA) mechanism for enhanced modality interaction. Merging blocks apply weighted pooling to downsample token sequences, while extension blocks use repeat up-sampling. Fast-StructText introduces new self-supervised tasks alongside MVLM: Graph-based Token Relation (GTR) for learning spatial relations, Sentence Order Prediction (SOP) for semantic continuity, and Text-Image Alignment to strengthen cross-modal grounding.
@@ 4.2.2 Joint-grained Frameworks
MGDoc [mgdoc] is the first multimodal, multi-granular pretrained framework designed to enhance inter-granular learning and cross-modal fusion through multi-granular and cross-modal attention mechanisms. It uses pretrained language encoders [sentencebert] and vision backbones [resnet] to extract features across different granularities—from word to page. These features are combined with positional embeddings [faststructext] and modality-type embeddings [structext] to form the final input representations. To capture hierarchical relations among pages, entities, and words, two attention biases are added: a binary hierarchical bias indicating inclusion relations, and a relation bias based on relative bounding box positions. Cross-modal attention is used to integrate information across modalities. Pretraining involves three tasks: Masked Text Modeling (MTM), Masked Vision Modeling (MVM), and Multi-Granularity Modeling. MTM and MVM mask and reconstruct multi-granular representations using mean absolute error, while token-entity linking is modeled via dot-product similarity to promote multi-granular alignment.
@@
WUKONG [wukong] utilizes fine-grained inputs (tokens with bounding boxes and document images) but enhances learning through coarse-grained self-supervised tasks. A Mask-RCNN backbone extracts visual features from textlines via RoIHead, while RoBERTa [wei2020robust] encodes token representations in two stages: the first six layers for initial textual features (augmented with layout features [layoutlmv2]), and the remaining layers for multimodal fusion with visual features. Four pretraining tasks are designed: Masked Language Modeling (MLM), Textline-Region Contrastive Learning (TRC), Masked Region Modeling (MRM), and Textline Grid Matching (TCM). TRC enhances entity-level cross-modal alignment via contrastive learning [yao2021filip]. MRM randomly masks 15% of textlines and predicts their visual embeddings [selfdoc]. TCM divides the image into grids and predicts the grid location for masked textlines, strengthening layout awareness. Each task’s loss is weighted with scaling parameters during training.
@@
GeoLayoutLM [geolayoutlm] is a sophisticated multimodal framework that distinctively incorporates geometric information through specialised pretraining tasks and the development of innovative relation heads. Inspired by the dual-stream structure of METER and SelfDoc [selfdoc], GeoLayoutLM features separate vision and text-layout modules coupled with interactive co-attention layers that enhance the integration of visual and textual data. The model introduces two advanced relation heads—the Coarse Relation Prediction (CRP) head and the Relation Feature Enhancement (RFE) head—which refine relation feature representation crucial for both pretraining and fine-tuning phases. The pretraining regimen includes tasks designed to understand geometric relationships, such as GeoPair, GeoMPair, and GeoTriplet, aiding the model in grasping the complex dynamics of document layouts. During fine-tuning, the model utilises pretrained parameters to optimise both semantic entity recognition and relation extraction tasks, employing a novel inference technique that enhances relation pair selection accuracy by focusing on the most probable relationships and minimizing variance among potential options.
4.3 Encoder-Decoder Pretrained Frameworks
@@
In addition to the above encoder-only frameworks, researchers have proposed encoder-decoder pretrained models that often approach tasks like Key Information Extraction (KIE) or Visual Question Answering (VQA) in a generative style. Addressing limitations of OCR-dependent frameworks, such as accumulated OCR errors and incorrect reading orders, OCR-free models have been introduced for end-to-end VRD-CU.
4.3.1 OCR-dependent Encoder-Decoder Frameworks
@@
TILT [tilt] is a T5-based transformer encoder-decoder architecture enhanced with a relative spatial bias in the self-attention mechanism to acquire fine-grained token representations. It incorporates encoding methods that apply relative sequence input bias and capture horizontal and vertical distance biases in the attention scores. A U-Net-based framework is applied to extract the fixed-size feature maps fed into the encoder together. They follow T5 pretraining strategies on RVL-CIDP dataset [rvlcdip] but use a salient span masking scheme adopted by [t5, guu2020retrieval]. As the first encoder-decoder framework, TILT requires off-the-shelf OCR tools to acquire the textual token sequence. To conduct VRD content understanding end-to-end, some OCR-free frameworks are proposed to solve the limitations of OCR-dependent models.
UDOP [udop] introduces an encoder-decoder framework for document understanding, built on a ViT-based model [vit] and inspired by LayoutLMv3 [layoutlmv3]. It enhances textual token embeddings by summing them with aligned image patch embeddings when the token’s bounding box center overlaps with a patch. TILT-style positional biases are used, while ID embeddings are omitted. The decoder consists of two cross-attending modules: a bidirectional Transformer for text-layout decoding and an MAE-style vision decoder. Pretraining includes both self-supervised tasks (e.g., masked text-layout and image reconstruction) and supervised tasks using benchmark datasets. Additional objectives include layout moulding (predicting token group positions) and visual text recognition (identifying text at specific locations). Supervised fine-tuning is performed on datasets for classification [rvlcdip], KIE [nda], VQA [visualmrc, docvqa], and layout analysis [publaynet].
@@
DocFormerv2 [docformerv2] is an encoder-decoder transformer architecture that uses multimodal (visual, textual and positional) to enhance the multimodal understanding and layout-aware language decoder to predict the predictions. The patched image pixels are fed into the convolutional and linear layers to get down-sampled patch embedding. The textual embeddings are acquired by linear projected token one-hot encoding. Both visual and textual embeddings are summed with the 2D-positional encoding [layoutlm] of Patches and linear project bbox embedding ($[x0,y0,x1,y1]$) [lilt], respectively. Two encoder-based Token to Line (T2L), Token to Grid (T2G), and one decoder-based self-supervised learning task, MLM, are proposed to enable multimodal feature interactive learning. T2L aims to improve the relative position understanding between tokens by predicting the number of textlines between two randomly selected tokens. For improving the layout and structure, understanding needs to split the image into $m\times n$ grids to predict the located grid number of each OCR token. For the decoder MLM, the spatial feature of each masked text token is masked as well, and the other setup follows T5 [t5].
@@
ViTLP [vitlp] presents an encoder-decoder framework for OCR and document understanding. It employs a ViT-based vision encoder to extract image patch representations, which are decoded autoregressively into text and layout sequences. A special “[LOC]” token encodes bounding box coordinates $[x_{0},y_{0},x_{1},y_{1}]$ to reduce layout sequence length. To guide generation, “[BOS]” and “[CONT]” tokens are used—representing a token pair t1, t2 as “[BOS], t1, [LOC], t2, [LOC]”. The decoder has hierarchical heads: the text head predicts the next token using the full sequence, while the layout head predicts bounding boxes only from “[LOC]” tokens. The “[CONT]” token enables variable-length generation by continuing decoding until “[END]”, based on the prefix token ratio.
4.3.2 OCR-free Pretrained Frameworks
@@
SeRum [serum] introduces an end-to-end OCR-free framework that decodes text based on local visual cues and spatial focus. A Swin Transformer extracts image patch features, which are refined through a Query Decoder to form vision-enhanced query representations, subsequently used by a Text Decoder for auto-regressive text generation. To better capture relevant regions, a Content-aware Token Merge module selects top-$K$ visual tokens based on their correlation with the query, while unselected background tokens enhance the foreground via attention. Three pretraining tasks are proposed: Query-to-Segmentation (Q2S), which predicts instance masks for text regions; Text-to-Segmentation (T2S), which segments based on the decoder’s output; and Segmentation-to-Text (S2T), which generates text from segmented features. These tasks jointly enhance SeRum’s ability to detect, align, and generate text directly from document images.
@@
StrucTextV2 [structextv2] is an end-to-end structure that uses image-only input to conduct several downstream tasks. It contains a CNN-based visual extractor with FPN strategies [lin2017feature] and follows ViT [vit] to get linear projected flattened patch-level representations. The patch token embeddings serve as the input to the Transformer encoder to enhance the contextually semantic representations. Then, the lightweight fusion network is applied to generate the final representations and fed into two branches during pretraining: made language Modeling (MLM) and Masked Image Modeling (MIM). Instead of using text inputs when MLM is used by other models [bert], a portion of the text regions are masked with RGB values $[255,255,255]$ randomly with a 2-layer MLP decoder to predict the masked token. MIM masks the rectangular text regions and predicts the RGB values of the missing pixels to improve the document representations. Except for the global average pooled FPN fused visual representations, the MLM-generated hidden state of each text region is concatenated and fed into a Fully Convolutional New York to get the regressed masked missing pixel values.
4.3.3 LLM-based Frameworks
With the rapid development of LLMs/MLLMs, many frameworks are adopted to tackle document-based tasks. Both open-source [qwen, llava] or close-source frameworks achieved promising performance on many benchmark datasets. This section introduces frameworks specifically proposed to address document related techniques with novel architectures. The trend of applying general domain LLMs/MLLMs in VRD-CU are discussed in Section 6.5.
DocLLM [docllm] proposes to address VRD-CU tasks by designing a disentangle spatial attention mechanism to capture alignments between text and layout modalities instead of integrating heavy visual backbones. The attention score between tokens $i$ and $j$ is computed as:
$A_{i,j}=Q^{t}_{i}(K^{t}_{j})^{\top}+\lambda_{t,s}Q^{t}_{i}(K^{s}_{j})^{\top}+\lambda_{s,t}Q^{s}_{i}(K^{t}_{j})^{\top}+\lambda_{s,s}Q^{s}_{i}(K^{s}_{j})^{\top}$ (13)
where $Q^{t}$, $K^{t}$ and $Q^{s}$, $K^{s}$ are projections from textual and spatial embeddings, respectively. Pretraining adopts an autoregressive block infilling objective to better handle irregular layouts, where coherent OCR-derived blocks are masked and predicted given both prefix and suffix context. DocLLM is then instruction-tuned on 16 datasets covering four tasks (VQA, KIE, etc) using natural prompts, enabling strong generalization without vision encoders.
LapDoc [lapdoc] enhances LLM-based VRD-CU framework by introducing a Layout-Aware Positional (LAP) embedding mechanism. It combines a learnable layout embedding $e_{l}$ with the standard sinusoidal token positional encoding $e_{p}$, forming a unified input representation $x=e_{t}+e_{p}+e_{l}$, where $e_{t}$ is the text embedding. The LAP encoder maps 2D bounding boxes into embeddings that capture spatial structure while preserving the autoregressive training scheme. Pretraining is conducted using standard language modeling on 240M document pages from the GLaMM corpus, integrating both text and layout features. Instruction tuning is performed on 9 datasets spanning four tasks—VQA, classification, KIE, and NLI—using task-specific prompt templates to guide model behavior, thereby enabling generalization across diverse document understanding tasks.
Some positional encoding strategies are proposed to enhance LLM/MLLM-based VRD-CU applications. LayTokenLLM introduces a layout token that summarizes the spatial configuration of each text segment using an attention-based layout tokenizer. This layout token is inserted alongside the text tokens and shares the same positional index as the first token in the segment, allowing the model to integrate layout cues without disrupting the sequential language modeling objective. By maintaining a full 100% T-Ratio (i.e., all positions contribute to text prediction), LayTokenLLM ensures efficient learning and inference, outperforming earlier layout-as-token baselines. In parallel, Group Position Embedding (GPE) proposes a complementary approach by assigning different spatial features—such as bounding box coordinates and reading order—to distinct attention head groups within rotary or sinusoidal positional encodings. This CoT-style decomposition enables the model to reason over multiple layout dimensions independently, enhancing spatial awareness without altering the model architecture. While LayTokenLLM focuses on embedding layout into the input sequence, GPE injects layout awareness through grouped attention, and both methods demonstrate strong performance across document benchmarks with minimal architectural overhead.
DocLayLLM incorporates explicit visual information by encoding image patches of document regions into external memory via a vision encoder (e.g., DiT), enabling stronger grounding in complex layouts. Crucially, it introduces Chain-of-Thought (CoT) prompting by fine-tuning on multi-step reasoning sequences, which guide the model to generate intermediate layout-aware explanations before predicting the final answer. This visual CoT-enhanced framework significantly boosts multi-hop document reasoning and better captures hierarchical and spatial relationships, outperforming prior text-only models on VQA, KIE, and NLI benchmarks.